Suppose you built all the rectangles possible from 48 square tiles. Which rectangle would have the largest perimeter? Which would have the smallest perimeter?
Answer by checkley77(12844)

2L+2W=PERIMETER.
L*W=48
1 BY 48=2*1+2*48=2+96=98 FOR THE LONGEST PERIMETER.
2 BY 24=2*2+2*24=4+48=52
3 BY 16=2*3+2*16=6+32=38
4 BY 12=2*4+2*12=8+24=32
6 BY 8=2*6+2*8=12+16=28 FOR THE SHORTEST PERIMETER.